OpenStudy (rockets31):

ok how would you write an equation in standard form when you have the points (-1,1) and (3,2)

OpenStudy (anonymous):

the standard form of an equation of a line is: y = mx+b the slope of a line, m, is (change in y/change in x) = (y2-y1)/(x2-x1) = (2-1)/(3--1) = 1/2 you can put the coordinates of one point into the equation now to solve for b (the y-intercept). y = mx+b 1 = (1/2)(-1) +b 1 = -1/2 + b 3/2 = b = y-intercept. so now, the standard form of an equation of a line is: y = mx+b so y = 1/2x+3/2